# Revealing Human Internal Attention Patterns from Gameplay Analysis for Reinforcement Learning

### v2

@@ Abstract
This study introduces a novel method for revealing human internal attention patterns from gameplay data alone, leveraging offline attention techniques from reinforcement learning (RL). We propose contextualized, task-relevant (CTR) attention networks, which generate attention maps from both human and RL agent gameplay in Atari environments. To evaluate whether the human CTR maps reveal internal attention, we validate our model by quantitative and qualitative comparison to the agent maps as well as to a temporally integrated overt attention (TIOA) model based on human eye-tracking data. Our results show that human CTR maps are more sparse than the agent ones and align better with the TIOA maps. Following a qualitative visual comparison we conclude that they likely capture patterns of internal attention. As a further application, we use these maps to guide RL agents, finding that human internal attention-guided agents achieve slightly improved and more stable learning compared to baselines. This work advances the understanding of human-agent attention differences and provides a new approach for extracting and validating internal attention from behavioral data.
@@ 1 Introduction
Figure 1: The goal of this study is to reveal human internal attention patterns solely from gameplay analysis: An attention model is proposed that extracts contextually task-relevant features from human and reinforcement learning (RL) agent gameplay. A temporally integrated gaze prediction model is used as a point of comparison to confirm that the attention model successfully reveals human internal attention.
@@
The frame problem [1] represents an ongoing challenge in artificial intelligence (AI) research: How does an agent determine what information is relevant in complex, ever-changing environments? Understanding how humans allocate attention in task-based scenarios can provide valuable insights into this problem to eventually design more intelligent, autonomous agents. By integrating research on human attention with visual attention-based reinforcement learning (RL), we can gain a deeper understanding of attention’s role in human task learning while also improving the design of intelligent, autonomous agents.
@@
As visual attention methods become more common in deep learning [2], several recent studies have explored incorporating attention mechanisms including human attention information into RL. These methods aim to either make an agent’s actions more interpretable, contributing to explainable AI, or improve learning efficiency and final performance. The attention approaches can be classified as offline-trainable and online-trainable. In offline-trainable attention methods such as [3, 4, 5, 6, 7, 8, 9, 10], an attention or saliency map is derived in a self-supervised way—either from the state (i.e., visual input) or from state-action-reward transitions (i.e., experience replay). This map is then used for pre-processing the input image or features by weighting or selection before processing in the policy network, or can even be applied after training for interpretation purposes too. In online methods such as [11, 12, 13, 14], attention is integrated directly into the learning process, for example, by being determined bymadditional attentive actions assigned to the agent, as in active vision reinforcement learning [11]. These attention mechanisms have been widely used in visual deep RL for Atari games but have also been applied to robotic navigation [15] and manipulation tasks [16].
@@
Apart from applying an agent’s own attention map, agents can also be augmented with human attention information. This has been shown to improve learning performance in imitation learning (IL), including behavioral cloning [17] and inverse RL [18, 19]. Human attention information can be obtained through recording eye-tracking data during gameplay [17], and predicted using gaze prediction models [20, 21]. There is also research on task-oriented human gaze and attention prediction. In [20], a top-down gaze prediction model is built using a dynamic Bayesian network to predict human gaze positions during gameplay. Yang et al. [21] used inverse reinforcement learning to model task-oriented attention in visual search paths on images. However, to fully understand human decision making in task-based scenarios, actual or predicted gaze data alone is not sufficient. This is because gaze data only reveals overt attention [22], i.e., the attention visible to an observer. Covert attention is the mental focus on something without necessarily looking at it—such as an object in the visual periphery or in memory. Because it involves no observable movements, it can not be measured by eye tracking and is challenging to model [20, 23, 24].
@@
Guo et al. [23] have done an in-depth comparison of human gaze prediction with RL agents’ attention (saliency map), leading to the insight that humans attend to fewer, less long-term objects. Also, agents tended to fail in situations where their attention strongly differed from the human attention. This study points out the limitations of using gaze data (overt attention) only to study human decision making and to build better human attention-augmented agents. If we had access to human internal attention maps—including both task-relevant overt and covert attention-they could be used to further improve intelligent, autonomous agent design. Without the need for additional brain activity data (such as in [24]), it remains to be explored whether task-relevant internal attention can be inferred from gameplay.
@@
In this paper, we define human internal attention as the combination of task-relevant overt attention (observable through gaze), covert attention (mental focus not accompanied by eye movements), and short-term visual memory, all corresponding to a given state. Together, these represent the full set of task-relevant objects that guide decision making. We argue that this internal attention is of particular importance for studying human decision making and for developing better human attention-augmented agents.
@@
The two main contributions of this study are the following: (1) We derive internal attention patterns from gameplay. The approach of this study is visualized in Fig. 1. (1a) We propose a contextualized, task-relevant (CTR) attention network trained hand-in-hand with a human action prediction network, inspired by offline attention methods in reinforcement learning (similar to visual attention in behavioral cloning [25]). Applied to the gameplay (experience replay) of humans and RL agents playing Atari games, this network helps to select features relevant to the current task and context at various sparsity levels. We first validate the CTR network using generated human and agent attention maps. (1b) We compare these attention maps against a custom temporally-integrated overt attention (TIOA) network, trained on eye-tracking data from the same human Atari gameplay. The TIOA serves as a point of comparison for covert attention, accounting for visual integration but not peripheral attention. Through this comparative analysis, we assess the CTR network’s ability to reveal internal attention patterns and provide insights that go beyond previous gaze-based studies. (2) We apply the derived internal attention maps for learning, guiding RL agents and evaluating their impact on learning performance. By leveraging these maps, we aim to transfer not only actions but also underlying attentional strategies, offering a complementary approach to imitation learning that may improve efficiency, robustness, and interpretability.
@@ 2 Methods
Figure 2: a) Contextualized task-relevant (CTR) attention network as well as temporally-integrated overt attention (TIOA) network architecture; b) Action predictor architecture; c) Autoencoder (AE) architecture; d) Method of creating a blended feature space $\mathcal{F}_{\mathrm{B}}$ from a target and source for the action predictor.
@@ 2.1 Study Overview and Datasets
For each game, and for both agent and human replay memory, we train one attention model (CTR attention) and one human gaze predictor (TIOA model). Training the attention network also involves an autoencoder and an action predictor. To prevent overreliance on elements apart from the main scene, interface elements and score indicators are deactivated. The neural network models are implemented and trained in PyTorch, with replay memory randomly split into 80% training and 20% validation transitions. All models are trained for at least 300 epochs until convergence. [^2] The code is made publicly available at: https://github.com/UThenrik/revealing-attention
@@ 2.2 Contextualized, Task-Relevant (CTR) Attention Network
based on the current state $\bm{s}\in\mathbb{R}^{4\times 84\times 84}$ and a sparsity controlling value $\lambda\in[0,1]$, parameterized by $\theta$. The state is represented as a stack of four game frames, each being a grayscale image of size $3\times 84\times 84$, downscaled from an RGB image of size $3\times 210\times 160$ as commonly done in Atari RL research. The state is first processed through a three-stage convolutional encoder, similar to the architecture in [6], reducing spatial resolution fourfold. The encoder weights are initialized from a pre-trained autoencoder, as illustrated in Fig. 2(b) and described in the next paragraph. After encoding, the features are contextualized using four serial dilated convolution layers with increasing dilation rates, concatenated with global maximum and average pooling channels, total spatial coordinates (similar to CoordConv [28]), and the expanded attention sparsity value $\lambda$. The final attention map is generated through three serial pointwise convolution layers, outputting a single-channel map of size $1\times 21\times 21$. Leaky ReLU activations are applied between layers, with a sigmoid activation at the output to ensure values remain in the $[0,1]$ range.
For training the CTR attention network, we use a pre-trained autoencoder (AE), as illustrated in Fig. 2(b). This AE is trained for each game separately on states from both agent and human replay memory and encodes the state into a feature tensor $\bm{\mathrm{F}}\in\mathbb{R}^{32\times 21\times 21}$ with 32 feature channels and spatial dimensions matching $\bm{\Psi}$. After the first two encoder layers we apply BlurPooling [29] with stride of 2 to ensure the feature space is shift-invariant while allowing for fourfold reduction of spatial resolution.
By applying the attention map at this encoded level, we achieve a balanced mix of spatial, feature-based, and object-based attention. For example, a single feature can preserve spatial location while also encoding an object (e.g., a car’s position in Freeway) and/or a higher-level representation (e.g., the car’s driving direction). The CTR attention network is trained alongside an action predictor, whose architecture is shown in Fig. 2(c). The action predictor estimates the action for the current target state based on a feature space:
@@
This blended space overlays the target feature tensor $\bm{\mathrm{F}}_{\mathrm{T}}$ and the source feature tensor $\bm{\mathrm{F}}_{\mathrm{S}}$, which is derived from a randomly sampled state. Here, $\odot$ and $\circ$ denote element-wise multiplication and exponentiation, respectively. The blending process for the feature spaces is visualized in Fig. 2(d).
@@
The sparsity-controlling factor $\lambda$ restricts the average activation of the attention map, specifying the maximum ratio of features that can be active. The loss term $\mathcal{L}_{\lambda,2}$ enforces this. While we assume that more features may improve action prediction by providing more information, we want to avoid unnecessary activations. The auxiliary loss $\mathcal{L}_{\lambda,1}$ also controlling minimum sparsity in early training is introducted for training stabilization.
@@
It is noted that the CTR attention network can be used plug-and-play with the self-supervised attention network from [6], which has been shown to improve the rate of convergence and performance of deep RL methods in Atari games, while additionally offering task-relevant and contextualized feature selection. This will therefore be investigated in Sec. 5. It could also be trained using inverse dynamics prediction, i.e. predicting an action $a_{k}$ from the current state $\bm{s}_{k}$ and the next $\bm{s}_{k+1}$. This approach would extract all features that are evidence of an action and could be more robust for applications in RL. Forward action prediction, equivalent to behavioral cloning, is used in this study as it leads the attention map to extract features that likely caused the player’s decision.
@@ 2.3 Temporally-Integrated Overt Attention (TIOA) Network
In order to analyze whether the CTR network can reveal human internal attention, a point of comparison is needed. Rather than using only the instantaneous gaze positions (eye-tracking data) from the Atari-HEAD dataset, an integration of recent gaze positions is preferable, as it more closely mimics the temporal integration of visual sensory data in humans [30]. This is relevant for comparison in accordance to our definition in Sec. 1 where human internal attention encompasses short-term visual memory as well. To further enhance generalization to unseen states (and those of the agent), we propose a temporally-integrated overt attention (TIOA) network. It features the same base architecture as the CTR attention network in Fig. 2(a) and therefore predicts temporally-integrated overt attention on the fourfold resolution-reduced feature space. Its output shall denote a probability distribution and therefore a softmax function is used after the final layer.
@@ 3 Evaluation I: Action Prediction under Attention Masking
Figure 4: Action predictor accuracies over unseen attention-masked states for human (CTR and TIOA) and agent (only CTR) gameplay at different attention activation rates/ sparsities.
@@
We observe that accuracies for the human actions are higher than in the original Atari-HEAD study [17], likely as we are predicting based on frame-stacked states.
In total, four main findings can be observed. Firstly, higher activation rates, i.e., more information leads to more accurate action prediction. This is the case for all games except Freeway. Secondly, the CTR attention maps are more efficient than the gaze data-based TIOA maps, achieving significantly higher accuracies with fewer activated features. This may be expected, as the CTR networks are optimized for this task. However it underlines the core motivation for this study as posed in Sec. 1, that the full internal attention including covert attention is required to better describe human decision making. The dashed lines in Fig. 4 indicate the minimum activation rate where 95% of the maximum action prediction accuracy is reached. Thirdly, human actions can be accurately predicted with less features than agent actions. This agrees with the finding of Guo et al. [23] which was based only on gaze data. Humans focus on less objects than agents trained for long-term future reward, even if we include potential objects of covert attention in the periphery beyond the gaze. Lastly, despite the expected greater gameplay variety in the human replay memory (due to the involvement of four different human subjects), the prediction accuracies are similar to those of a replay memory of a single trained agent.
The findings suggest that the CTR network correctly identifies contextually task-relevant features. Whether these maps plausibly resemble human internal attention patterns will be further analyzed quantitatively and qualitatively in Sec. 4.
4 Evaluation II: Analysis Revealed Attention Maps

### v3

# Revealing Human Attention Patterns from Gameplay Analysis for Reinforcement Learning

@@ Abstract
This study introduces a novel method for revealing human internal attention patterns (decision-relevant attention) from gameplay data alone, leveraging offline attention techniques from reinforcement learning (RL). We propose contextualized, task-relevant (CTR) attention networks, which generate attention maps from both human and RL agent gameplay in Atari environments. To evaluate whether the human CTR maps reveal internal attention patterns, we validate our model by quantitative and qualitative comparison to the agent maps as well as to a temporally integrated overt attention (TIOA) model based on human eye-tracking data. Our results show that human CTR maps are more sparse than the agent ones and align better with the TIOA maps. Following a qualitative visual comparison we conclude that they likely capture patterns of internal attention. As a further application, we use these maps to guide RL agents, finding that human attention-guided agents achieve slightly improved and more stable learning compared to baselines, and significantly outperform TIOA-based agents. This work advances the understanding of human-agent attention differences and provides a new approach for extracting and validating internal attention patterns from behavioral data.
@@ 1 Introduction
Figure 1: The goal of this study is to reveal human internal attention patterns (decision-relevant attention) solely from gameplay analysis: An attention model is proposed that extracts contextually task-relevant features from human and reinforcement learning (RL) agent gameplay. A temporally integrated gaze prediction model is used as a point of comparison to confirm that the attention model successfully reveals human internal attention patterns.
@@
The frame problem [mccarthy1981some] represents an ongoing challenge in artificial intelligence (AI) research: How does an agent determine what information is relevant in complex, ever-changing environments? Understanding how humans allocate attention in task-based scenarios can provide valuable insights into this problem to eventually design more intelligent, autonomous agents. By integrating research on human attention with visual attention-based reinforcement learning (RL), we can gain a deeper understanding of attention’s role in human task learning while also improving the design of intelligent, autonomous agents.
@@
As visual attention methods become more common in deep learning [hassanin2024visual], several recent studies have explored incorporating attention mechanisms including human attention information into RL. These methods aim to either make an agent’s actions more interpretable, contributing to explainable AI, or improve learning efficiency and final performance. The attention approaches can be classified as offline-trainable and online-trainable. In offline-trainable attention methods such as [greydanus2018visualizing, manchin2019reinforcement, gopalakrishnan2020unsupervised, wu2021self, shi2020self, nikulin2019free, ma2025don, chen2024focus], an attention or saliency map is derived in a self-supervised way—either from the state (i.e., visual input) or from state-action-reward transitions (i.e., experience replay). This map is then used for pre-processing the input image or features by weighting or selection before processing in the policy network, or can even be applied after training for interpretation purposes too. In online methods such as [shang2023active, mnih2014recurrent, kulkarni2019unsupervised, mott2019towards], attention is integrated directly into the learning process, for example, by additional attentive actions assigned to the agent, as in active vision reinforcement learning [shang2023active]. These attention mechanisms have been widely used in visual deep RL for Atari games but have also been applied to robotic navigation [salter2021attention] and manipulation tasks [james2022coarse].
@@
Apart from applying an agent’s own attention map, agents can also be augmented with human attention information. Zhang et al. [zhang2019leveraging] provide a comprehensive survey of methods for leveraging human guidance in deep RL, categorizing attention-based learning as one approach for agents to learn from humans. This has been shown to improve learning performance in imitation learning (IL), including behavioral cloning [zhang2020atari] and inverse RL [zhang2018agil, saran2021efficiently]. Human attention information can be obtained through recording eye-tracking data during gameplay [zhang2020atari], and predicted using gaze prediction models [borji2014what, yang2020predicting]. However, as Zhang et al. [zhang2019leveraging] note, existing work in this area has primarily focused on using gaze (overt attention) as the sole attention signal. There is also research on task-oriented human gaze and attention prediction. In [borji2014what], a top-down gaze prediction model is built using a dynamic Bayesian network to predict human gaze positions during gameplay. Yang et al. [yang2020predicting] used inverse reinforcement learning to model task-oriented attention in visual search paths on images. However, to fully understand human decision making in task-based scenarios, actual or predicted gaze data alone is not sufficient. This is because gaze data only reveals overt attention [posner1980orienting], i.e., the attention visible to an observer. Covert attention is the mental focus on something without necessarily looking at it—such as an object in the visual periphery or in memory. Because it involves no observable movements, it can not be measured by eye tracking and is challenging to model [borji2014what, guo2021machine, leong2017dynamic].
@@
Guo et al. [guo2021machine] conducted an in-depth comparison of human gaze prediction with RL agents’ perturbation-based attention (saliency map), showing that agent attention becomes more similar to human gaze when the discount factor is reduced from the typical 0.99 (long-term reward focus), and that agents fail more often when their attention diverges from human gaze. However, their analysis relied exclusively on gaze (overt attention) and therefore did not account for covert attention or short-term memory. This study points out the limitations of using gaze data (overt attention) only to study human decision making and to build better human attention-augmented agents. If we had access to human attention patterns—including both decision-relevant overt and covert attention-they could be used to further improve intelligent, autonomous agent design. Without the need for additional brain activity data (such as in [leong2017dynamic]), it remains to be explored whether these internal attention patterns can be inferred from gameplay.
@@
In this paper, we define human internal attention patterns (or equivalently, decision-relevant attention patterns) as generalized patterns reflecting the combination of overt attention (observable through gaze), covert attention (mental focus not accompanied by eye movements), and short-term visual memory, all corresponding to a given state. Together, these represent the full set of features a human typically uses to make decisions. We use the term “internal” to distinguish this from agent-optimized task-relevant attention, as human internal attention is specific to human decision-making and independent of whether those decisions are optimal for task performance. We argue that these internal attention patterns are of particular importance for studying human decision making and for developing better human attention-augmented agents. To extract these patterns from gameplay, we couple action prediction with attention extraction. While similar methods have been used to explain agent policies (e.g., Chen et al. [chen2024focus], Shi et al. [shi2022temporal], Greydanus et al. [greydanus2018visualizing]), our focus is on deriving attention maps that explain human decision-making.
@@
The two main contributions of this study are the following: (1) We derive internal attention patterns from gameplay. The approach of this study is visualized in Fig. 1. (1a) We propose a contextualized, task-relevant (CTR) attention network trained hand-in-hand with an action prediction network, inspired by offline attention methods in reinforcement learning (similar to visual attention in behavioral cloning [mohandoss2019visual]). Applied to the gameplay (experience replay) of humans and RL agents playing Atari games, this network helps to select features relevant to the current state at various sparsity levels. We first validate the CTR network using generated human and agent attention maps. (1b) We compare these attention maps against a custom temporally-integrated overt attention (TIOA) network, trained on eye-tracking data from the same human Atari gameplay. The TIOA serves as a point of comparison for covert attention, accounting for visual integration but not peripheral attention. Through this comparative analysis, we assess the CTR network’s ability to reveal internal attention patterns and provide insights that go beyond previous gaze-based studies. (2) We apply the derived internal attention patterns for learning, guiding RL agents and evaluating their impact on learning performance. We compare CTR-masked agents against TIOA-masked agents, demonstrating that CTR attention can offer additional value beyond overt attention alone for RL applications. By leveraging these maps, we aim to transfer not only actions but also underlying attentional strategies, offering a complementary approach to imitation learning that may improve efficiency, robustness, and interpretability.
@@ 2 Methods
e)Refer to caption
Figure 2: a) Contextualized task-relevant (CTR) attention network as well as temporally-integrated overt attention (TIOA) network architecture; b) Autoencoder (AE) architecture; c) Action predictor architecture; d) Method of creating a blended feature space $\mathcal{F}_{\mathrm{B}}$ from a target and source for the action predictor; e) Exmple attention map and ablation study: Example frame and attention map (left), and log10 histogram of activations (right) over different sparsity control factors $\lambda$ for the human CTR attention network with and without feature blending. Feature blending drives the activation distribution toward binarization (values near 0 or 1).
@@ 2.1 Study Overview and Datasets
Although four subjects were recorded for each game, the available data per individual is deemed insufficient for robust subject-specific analysis. We therefore aggregate the gameplay data from all subjects to focus on more general differences between human and agent attention patterns. Consequently, an analysis of individual differences among subjects was not performed in this study. For each game, and for both agent and human replay memory, we train one attention model (CTR attention) and one human gaze predictor (TIOA model). Training the attention network also involves an autoencoder and an action predictor. To prevent overreliance on elements apart from the main scene, interface elements and score indicators are deactivated. The neural network models are implemented and trained in PyTorch, with replay memory randomly split into 80% training and 20% validation transitions. All models are trained for at least 300 epochs until convergence. [^2] The code is made publicly available at: https://github.com/UThenrik/revealing-attention
@@ 2.2 Contextualized, Task-Relevant (CTR) Attention Network
based on the current state $\bm{s}\in\mathbb{R}^{4\times 84\times 84}$ and a sparsity controlling value $\lambda\in[0,1]$, parameterized by $\theta$. The state is represented as a stack of four game frames, each being a grayscale image of size $3\times 84\times 84$, downscaled from an RGB image of size $3\times 210\times 160$ as commonly done in Atari RL research. The state is first processed through a three-stage convolutional encoder, similar to the architecture in [wu2021self], reducing spatial resolution fourfold. The encoder weights are initialized from a pre-trained autoencoder, as illustrated in Fig. 2(b) and described in the next paragraph. After encoding, the features are contextualized using four serial dilated convolution layers with increasing dilation rates, concatenated with global maximum and average pooling channels, total spatial coordinates (similar to CoordConv [liu2018intriguing]), and the expanded attention sparsity value $\lambda$. The final attention map is generated through three serial pointwise convolution layers, outputting a single-channel map of size $1\times 21\times 21$. Leaky ReLU activations are applied between layers, with a sigmoid activation at the output to ensure values remain in the $[0,1]$ range.
For training the CTR attention network, we use a pre-trained autoencoder (AE), as illustrated in Fig. 2(b). This AE is trained for each game separately on states from both agent and human replay memory and encodes the state into a feature tensor $\bm{\mathrm{F}}\in\mathbb{R}^{32\times 21\times 21}$ with 32 feature channels and spatial dimensions matching $\bm{\Psi}$. After the first two encoder layers we apply BlurPooling [pmlr-v97-zhang19a] with stride of 2 to ensure the feature space is shift-invariant while allowing for fourfold reduction of spatial resolution.
By applying the attention map at this encoded level, we achieve a balanced mix of spatial, feature-based, and object-based attention. For example, a single feature can preserve spatial location while also encoding an object (e.g., a car’s position in Freeway) and/or a higher-level representation (e.g., the car’s driving direction). The CTR attention network is trained alongside an action predictor, whose architecture is shown in Fig. 2(c). Crucially, we architecturally separate the attention network from the action predictor: the CTR network produces masks that are applied before the action predictor. This design ensures that extracted attention maps reflect genuine decision-relevant features rather than being artifacts of the action prediction optimization process itself. The action predictor estimates the action for the current target state based on a feature space:
@@
This blended space overlays the target feature tensor $\bm{\mathrm{F}}_{\mathrm{T}}$ and the source feature tensor $\bm{\mathrm{F}}_{\mathrm{S}}$, which is derived from a randomly sampled state. Here, $\odot$ and $\circ$ denote element-wise multiplication and exponentiation, respectively. The blending process for the feature spaces is visualized in Fig. 2(d). This feature blending mechanism drives the CTR attention network toward binarized attention maps. By requiring the network to distinguish between target and source features while maintaining action prediction accuracy, attention values are encouraged to converge toward either 0 (irrelevant) or 1 (relevant). An ablation study in Fig. 2(e) shows that without feature blending, the network produces less binarized activation distributions.
@@
The sparsity-controlling factor $\lambda$ restricts the average activation of the attention map, specifying the maximum ratio of features that can be active. The loss term $\mathcal{L}_{\lambda,2}$ enforces this. While we assume that more features may improve action prediction by providing more information, we want to avoid unnecessary activations. The auxiliary loss $\mathcal{L}_{\lambda,1}$ also controlling minimum sparsity in early training is introducted for training stabilization. The sparsity factor $\lambda$ is implemented as a controllable network input, enabling post-training adjustment across multiple sparsity levels essential for our comparative analysis in Sec. 3. Without regularization, attention maps converge to full activation to maximize information, defeating the purpose of selective attention.
@@
It is noted that the CTR attention network can be used plug-and-play with the self-supervised attention network from [wu2021self], which has been shown to improve the rate of convergence and performance of deep RL methods in Atari games, while additionally offering task-relevant and contextualized feature selection. This will therefore be investigated in Sec. 5. It could also be trained using inverse dynamics prediction, i.e. predicting an action $a_{k}$ from the current state $\bm{s}_{k}$ and the next $\bm{s}_{k+1}$. This approach would extract all features that are evidence of an action and could be more robust for applications in RL. However, we use forward action prediction as it directly extracts features that preceded and likely influenced the player’s decision, aligning with our goal to reveal decision-relevant attention patterns.
@@ 2.3 Temporally-Integrated Overt Attention (TIOA) Network
In order to analyze whether the CTR network can reveal human internal attention, a point of comparison is needed. Rather than using only the instantaneous gaze positions (eye-tracking data) from the Atari-HEAD dataset, an integration of recent gaze positions is preferable, as it more closely mimics the temporal integration of visual sensory data in humans [fairhall2014temporal]. This is relevant for comparison in accordance to our definition in Sec. 1 where human internal attention encompasses short-term visual memory as well. To further enhance generalization to unseen states (and those of the agent), we propose a temporally-integrated overt attention (TIOA) network. It features the same base architecture as the CTR attention network in Fig. 2(a) and therefore predicts temporally-integrated overt attention on the fourfold resolution-reduced feature space. Its output shall denote a probability distribution and therefore a softmax function is used after the final layer.
@@ 3 Evaluation I: Action Prediction under Attention Masking
Figure 4: Action predictor accuracies over unseen attention-masked states for human (CTR and TIOA) and agent (only CTR) gameplay at different attention activation rates/ sparsities. Dashed lines indicate the minimum activation rate where 97% of the maximum action prediction accuracy is reached.
Figure 5: Relative percentage decrease in action prediction accuracy from maximum comparing human CTR vs. agent CTR across different sparsity levels. Mean and 95% confidence interval across six games. Agent attention shows approximately twice the relative accuracy decrease compared to human attention, supporting that agents have more distributed attention.
@@
We observe that accuracies for the human actions are higher than in the original Atari-HEAD study [zhang2020atari], likely as we are predicting based on frame-stacked states.
In total, four main findings can be observed. Firstly, higher activation rates, i.e., more information leads to more accurate action prediction. This is the case for all games except Freeway. Secondly, the CTR attention maps are significantly more efficient than the gaze data-based TIOA maps, achieving higher action prediction accuracies with fewer activated features. This finding is supported by one-sided t-tests; for example, at a 16% activation rate (as used in the RL experiments in Sec. 5), where each game’s mean accuracy is treated as an independent sample, we observe that action prediction accuracy under the human CTR masks is significantly higher than under the TIOA masks ($p=0.0093$). This may be expected, as the CTR networks are optimized for this task. However it underlines the core motivation for this study as posed in Sec. 1, that the full internal attention including covert attention is required to better describe human decision making. The dashed lines in Fig. 4 indicate the minimum activation rate where 97% of the maximum action prediction accuracy is reached. Thirdly, human actions can be accurately predicted with fewer features than agent actions, consistent with the conclusion of Guo et al. [guo2021machine] from an analysis of human gaze data. Our results extend this finding to decision-relevant attention patterns extracted via behavioral cloning, capturing internal attention beyond the gaze peak. Fig. 5 quantifies this difference: agent attention shows approximately twice the relative accuracy decrease from maximum compared to human attention, though with large confidence intervals reflecting game-dependent variation. While this supports Guo et al.’s finding, direct quantitative comparison is limited as their work compared agent saliency maps with human gaze peaks (single concentrated regions), whereas our CTR approach captures broader internal attention including covert attention and peripheral awareness, suggesting the human-agent attention breadth difference may be less pronounced as it appears from gaze-focused analyses. Lastly, despite the expected greater gameplay variety in the human replay memory (due to the involvement of four different human subjects), the prediction accuracies vary by game but are generally similar to those of a replay memory of a single trained agent.
The findings suggest that the CTR network correctly identifies contextually decision-relevant features. Whether these maps plausibly resemble human internal attention patterns will be further analyzed quantitatively and qualitatively in Sec. 4.
4 Evaluation II: Analysis of Revealed Attention Maps